OpenStudy (anonymous):

Please check my answers I got A. for both! @ganeshie8

OpenStudy (help!!!!):

what was ur answer for the last one?

OpenStudy (anonymous):

@satellite73 Ahh.. Glad you decided to join us :) @HELP!!!! A for BOTH questions.

OpenStudy (anonymous):

i have no damned idea

OpenStudy (anonymous):

if you say what you did maybe i can understand it

OpenStudy (anonymous):

@HELP!!!! you know how to do these ?

OpenStudy (anonymous):

this is what i get for the last one http://www.wolframalpha.com/input/?i=2*23*72%2B2*23*84.25

OpenStudy (anonymous):

which fortunately is actually an answer, namely B

OpenStudy (help!!!!):

yeah on the last one i messed up lol

OpenStudy (help!!!!):

basically its the surface area

OpenStudy (help!!!!):

last one I have 8843.5

OpenStudy (anonymous):

i messed the last one up forgot the bottom

OpenStudy (anonymous):

man i must really be messing up

OpenStudy (anonymous):

lats one i get A like you do!

OpenStudy (anonymous):

Otay. For the second question: Two of the walls will have an area of 72⋅23= 1656 ; 1656 ⋅ 2 walls= 3312 the other two will have an area of 23⋅84.25= 1937.75 ; 1937.75⋅ 2 walls = 3875.5 and the floor will have an area of 72⋅ 84.25 = 6066 Add the 3 answers you got! 3312 + 3875.5 + 6066= 13253.5 , thus how I got option "A."

OpenStudy (anonymous):

that is what you get for asking for help lol lot of confusion but i think it is A and here is my computation http://www.wolframalpha.com/input/?i=2*23*72%2B2*23*84.25%2B72*84.25

OpenStudy (anonymous):

I just gave you the computation! -.-

OpenStudy (anonymous):

yeah it is the same as mine i got A too

OpenStudy (anonymous):

yay , no more confusion? -.- Also, did any of you figure out the first one?

OpenStudy (anonymous):

@HELP!!!!

OpenStudy (anonymous):

you said A i have no idea first you need to find that missing dimension from the box

OpenStudy (anonymous):

When in doubt call @ganeshie8 ...I don't think he's going to come though, he's usually not on and this time.

OpenStudy (anonymous):

at* Yay, he's here ! :DD

OpenStudy (anonymous):

ok i screwed up that is what you are looking for the surface area of the prism not the pyramid now i have to start all over

OpenStudy (anonymous):

doesnt the formula for the surface area of a prism have a square root in it?

OpenStudy (anonymous):

i get \(202\) rounded, i guess you go with \(205\)

ganeshie8 (ganeshie8):

for the pyramid, subtracting base area(6x5=30) from surface area gives the lateral area : `lateral area of pyramid = 107 - 30 = 77` multiplying it by 2 gives the lateral area of prism `lateral area of prism = 77*2 = 154` add this to the area of bases

OpenStudy (anonymous):

Thank you, that's what I got. :)